Take another endomesta student mu e s ste de tache Theme a) PE PO )
Answer #1

n=8 , p=112 ieve every areshen had two choiced so probability of corret apowessis ille X: Number of correct answer x » Binomi③ P(X33) = 1- P(x<3) =o) + P(x=1 &P(x=2)+ P.(x=3) = 1- [ 8c (0.53 © Coug& # 86, COSS COU5T+8c cart young (1x140.00394 (8 to 5